An object that follows Hooke’s Law is hanging 2 m above the ground with a mass of 1.5 kg and a spring constant of 1.2 N/m. If the object is at rest when stretched 0.25 m, what is its total mechanical energy?
1
Expert's answer
2021-08-02T08:41:48-0400

The total mechanical energy consists (in this case) of gravitational and elastic potential energies. The gravitational one is given as follows:


"E_g = mgh"

where "m = 1.5kg" is the mass of the object, "h = 2m" is the height above the ground, "g = 9.81m\/s^2" is the gravitational acceleration.

The elastic one:


"E_e = \dfrac{kx^2}{2}"

where "k = 1.2N\/m" is the spring constant, "x = 0.25m" is the elongatin of the spring.

Thus, obtain:


"E = E_g + E_e = 1.5\cdot 2\cdot 9.81 + \dfrac{1.2\cdot 0.25^2}{2} \approx 29J"

Answer. 29 J.